OpenStudy (goformit100):

Find the remainder when 2^1990 is divided by 1990.

OpenStudy (goformit100):

@mayankdevnani

Parth (parthkohli):

Mod arithmetic :') @terenzreignz

terenzreignz (terenzreignz):

Why me? :/

Parth (parthkohli):

Because you.

OpenStudy (goformit100):

In this question How to square to so much power ?

terenzreignz (terenzreignz):

Might have to resort to totients..... @ParthKohli ?

OpenStudy (goformit100):

Sir if i use the exponent rule will it work here ?

Parth (parthkohli):

Ah! Euler's Theorem!

OpenStudy (anonymous):

factor 1990 first

terenzreignz (terenzreignz):

Time to doodle... \[\large 2^{1990}=4^{995}\]

OpenStudy (goformit100):

How to factor it ?

OpenStudy (anonymous):

how to factor 1990?

terenzreignz (terenzreignz):

What is 4^5? \[\Large = 1024^{199}\]

OpenStudy (goformit100):

Yes @satellite73

OpenStudy (anonymous):

try \(2\times 5\times 199\)

OpenStudy (mayankdevnani):

1990 = 10*199 = 2* 5* 199

terenzreignz (terenzreignz):

\[\large 1024^{199}=1024\cdot 1024^{198}=1024\cdot 2048^{99}\]

OpenStudy (goformit100):

try \(2\times 5\times 199\) means ?

terenzreignz (terenzreignz):

Now let's start working some "mod magic" and reduce the bases at mod 1990 \[\Large =_{(mod \ 1990)} \ \ 1024\cdot 58^{99}\]

OpenStudy (mayankdevnani):